cons 3D SmartArt Chart My Add-ios - Comment Video Add-ins Text Symbols Nat iadiavor Company usoa a normal cog systas with a single Company uses a normal costing systom with a single MOH cost pool and MH as the Budgeted manufacturing overhead IMOH) Overhoad allocation base Budgeted machine-hours (MH) Manufacturing overhead (MOH) incurred Actual machine-hours (MH machine hou 75,000 as follows: 2019 End-of-Year Cost of Goods Sold 60,000 11,000 4,000 1,250,000 750,000 Work-in-Process Control 1. Compute the budgeted manufacturing overhead rate for 2019 2. Compute the under- or ovorallocated monufacturing overhead of Nat Radiator in 2019. Dispose of this under- or overallocated emount using: a. Write-off to Cost of Goods Sold b. Proration based on ending balances (before proration) in Work-in-Process Control, Finished Goods Control, and Cost of Goods Sold. c. Proration based on the allocated overhe Work-in-Process Control, Finishod Goods Control, and Cost of Goods Sold. 3. Which method do you prefer in requirement 27 Explain.
Answer #1
Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.
Requirement 1
Budgeted Manufacturing ovehread Rate:
Budgeted Manufacturing Overhead A $           4,800,000
Divided by: Budgeted Machine Hours B                    80,000 Hours
Budgeted Manufacturing ovehread Rate A/B $                       60
Requirement 2
Actual Overhead $           4,900,000
Applied Overheads (75,000 Hours*$ 60) $           4,500,000
Under applied overheads $              400,000
a. Write off to Cost of Goods Sold
Account Balance Write-off   After proration
A B C D=B+C
Work in Process $                            750,000 $                        -   $            750,000
Finished Goods $                         1,250,000 $                        -   $         1,250,000
Cost of Goods Sold $                         8,000,000 $              400,000 $         8,400,000
Total $                       10,000,000 $              400,000 $       10,400,000
b. Write off
Account Balance Ratio Write-off   After proration
A B C D=400000*C E=B+D
Work in Process $                            750,000 7.5% $              30,000 $           780,000
Finished Goods $                         1,250,000 12.5% $              50,000 $       1,300,000
Cost of Goods Sold $                         8,000,000 80.0% $            320,000 $       8,320,000
Total $                       10,000,000 100.0% $            400,000 $     10,400,000
c. Write off
Account Balance Balance included Ratio Write-off   After proration
A B Working below C D=400000*C E=B+D
Work in Process $                            750,000 $              240,000 5.3% $             21,333 $           771,333
Finished Goods $                         1,250,000 $              660,000 14.7% $             58,667 $       1,308,667
Cost of Goods Sold $                         8,000,000 $           3,600,000 80.0% $           320,000 $       8,320,000
Total $                       10,000,000 $           4,500,000 100.0% $           400,000 $     10,400,000
Working:
Work in Process 4000 Machine Hours*60
Finished Goods 11000 Machine Hours*60
Cost of Goods Sold 60000 Machine Hours*60
Requirement 3
Alternative C is prefereable as balances in this alternative is same which would have been in case of  
actual indirect cost usage